I'm making a book about the Pythagorean Theorem in my geometry class and one of the requirements is to find out why a squared + b squared= c squared is incorrect. I've done research on the theorem and still can't figure out why this is true. Can you please help me?
Answer by richwmiller(9135)

I suppose it all depends on what you think a^2+b^2=c^2 means.
If the legs of the right triangle are a and b and the hypotenuse is c then a^2+b^2=c^2.
Someone is pulling your leg.